In studies for a​ medication,

14

percent of patients gained weight as a side effect. Suppose

777

patients are randomly selected. Use the normal approximation to the binomial to approximate the probability that

​(a) exactly

109

patients will gain weight as a side effect.

​(b) no more than

109

patients will gain weight as a side effect.

​(c) at least

125

patients will gain weight as a side effect. What does this result​ suggest?
Answer #1

Mean = n * p = 777 * 0.14 = 108.78

Standard deviation = sqrt [ np(1-p) ]

= sqrt [ 777 * 0.14 ( 1 - 0.14) ]

= 9.6722

Using normal approximation

P(X < x) = P(Z < x - Mean / SD)

a)

With continuity correction,

P(X = 109) = P(108.5 < X < 109.5)

= P(X < 109.5) - P(X < 108.5)

= P(Z < (109.5 - 108.78) / 9.6722) - P(Z < (108.5 - 108.78) / 9.6722)

= P( Z < 0.07) - P(Z < -0.03)

= 0.5279 - 0.4880

= 0.0399

b)

P(X <= 109) = P(X < 109.5)

= P(Z < (109.5 - 108.78) / 9.6722)

= P(Z < 0.07)

= 0.5279

c)

P(X >= 125) = P(X > 124.5)

= P(Z > (124.5 - 108.78) / 9.6722)

= P(Z > 1.63)

= 0.0516